OpenStudy (pinkmilk234567):

6. f(x) = x3 + 1; {-2, -1, 3 i need help asap

OpenStudy (cas_fangirl_14):

what are you trying to find/answer?

OpenStudy (pinkmilk234567):

yes

OpenStudy (pinkmilk234567):

find the range of each function for the given domain

OpenStudy (kittiwitti1):

I'm not sure I understand the format of your question. What is {-2, 1, 3 ?

OpenStudy (pinkmilk234567):

idk

OpenStudy (kittiwitti1):

I assume your function equation is \(f(x)=x^{3}+1\). But what domains are listed by the problem?

OpenStudy (cas_fangirl_14):

Can you take a screenshot or picture of the question?

OpenStudy (pinkmilk234567):

6. f(x) = x3 + 1; {-2, -1, 3

OpenStudy (kittiwitti1):

x^3? And again... what is "{-2, -1, 3" ?

OpenStudy (pinkmilk234567):

idk

OpenStudy (kittiwitti1):

Do you have a physical form of the problem that can be copied virtually?

OpenStudy (kittiwitti1):

Oh, okay. I see it now... did you mean this? \(f(x) = x^{3} + 1; \left\{ -2, -1, 3 \right\}\)

OpenStudy (pinkmilk234567):

yes

Directrix (directrix):

I can't read the .doc but here is what I think the task it. Take the function f(x) = x^3 + 1 Evaluate f(x) = x^3 + 1 three times, once for each value of x in its domain of {-2, -1, 3}. Crank out f(-2) and then f(-1) and then f(3). Those 3 values you get will be the range of the function.

OpenStudy (kittiwitti1):

... I feel dumb now ... thanks @Directrix for clarifying though lol

Directrix (directrix):

@kittiwitti1 Back to you.

Directrix (directrix):

Don't feel dumb. The reason I had an idea of how to do it is that I saw a similar problem last week on OS.

OpenStudy (pinkmilk234567):

so whats the answer?

OpenStudy (kittiwitti1):

Ah, I see.... Still, thank you for helping me understand (: @Pinkmilk234567 try what he said, and tell me what you got, alright?

OpenStudy (pinkmilk234567):

ok

OpenStudy (pinkmilk234567):

i dont understand

OpenStudy (pinkmilk234567):

@kittiwitti1

OpenStudy (kittiwitti1):

He's saying put each of the domain values into the equation. Domain values are \(x\)-values, so you replace the \(x\) variable with each respective domain value.

OpenStudy (pinkmilk234567):

so could you do it for me @kittiwitti1

OpenStudy (kittiwitti1):

Sorry, no; direct answers/help isn't allowed on this site. It doesn't facilitate good learning/study habits.

OpenStudy (kittiwitti1):

But I will give you a hint: the values \(\left\{-2,1,3\right\}\) can be viewed as \(D:\left\{x_{1},x_{2},x_{3}\right\}\) If you put these into the function \(y=x^{3}+1\) then you would get the following functions:\[y_{1}=(x_{1})^{3}+1\]\[y_{2}=(x_{2})^{3}+1\]\[y_{3}=(x_{3})^{3}+1\]By the way, \(f(x)\) is another term for \(y\).

OpenStudy (kittiwitti1):

*Basically, you are getting three ranges (\(y\)-values).

Directrix (directrix):

The OP left us.

OpenStudy (kittiwitti1):

Actually, it's -1, I made a typo the second time around typing it @Directrix Sorry about that!

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\)\(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (kittiwitti1):

\(\text{LOL, good job}\) @likeabossssssss P:

likeabossssssss (likeabossssssss):

thx @kittiwitti1